A 2.0-g particle moving at 6.2 m/s makes a perfectly elastic head-on collision with a resting 1.0-g object.



(a) Find the speed of each particle after the collision.

2.0 g particle ___    m/s

1.0 g particle ___   m/s



(b) Find the speed of each particle after the collision if the stationary particle has a mass of 10 g.

2.0 g particle  ___   m/s

10.0 g particle ___   m/s



(c) Find the final kinetic energy of the incident 2.0-g particle in the situations described in parts (a) and (b).



KE in part (a) ___  J

KE in part (b) ___    J





1
Expert's answer
2021-10-05T15:44:20-0400

a)

"V_1 = (\frac{m_1-m_2}{m_1+m_2})v_1 + (\frac{2m_2}{m_1+m_2})v_2 \\

V_2 = (\frac{m_1-m_2}{m_1+m_2})v_2 + (\frac{2m_1}{m_1+m_2})v_1 \\

m_1=2g \\

m_2 = 1g \\

v_1 = 6.2 \;m/s \\

v_2 =0 \\

V_1 = (\frac{2-1}{2+1})6.2 +0 = 2.07 \;m/s \\

V_2 = 0 + \frac{2 \times 2}{2+1}6.2 = 8.27 \;m/s"

b)

"m_1=2g \\

m_2= 10g \\

v_1 = 6.2 \;m/s \\

v_2 = 0 \\

V_1 = (\frac{2-10}{2+10})6.2+0 = -4.13 \;m/s \\

V_2 = 0 +(\frac{2 \times 2}{2+10})6.2 = 2.07 \;m/s"

c)

"KE_a = \frac{1}{2}m_1v_1^2 \\

= \frac{1}{2} \times 2 \times 10^{-3} \times (2.07)^2 = 4.37 \times 10^{-3} \;J \\

KE_b = \frac{1}{2}m_1v_1^2 \\

= \frac{1}{2} \times 2 \times 10^{-3} \times (-4.13)^2 = 17.1 \times 10^{-3} \;J"